calculate GDP and National income: transfer Payment=80, govt. purchase=250, net indirect taxes=40, proprietrs income=160,wages & salaries=480,personal consump. exp.=500, capital cons. allowance=100, export-imp=50,corporate profit=120,interst &invest income=50,gross invest=150, net invest from non residents=60
1
Expert's answer
2018-05-10T09:59:08-0400
Calculate GDP and National income: transfer Payment = 80, govt. purchase = 250, net indirect taxes = 40, proprietors income = 160, wages & salaries = 480, personal consump. exp. = 500, capital cons. allowance = 100, export-imp = 50, corporate profit = 120, interest &invest income = 50, gross invest = 150, net invest from non residents = 60.
GDP = C + I + G + NX = 500 + (150 - 60) + 250 + 50 = 890.
NI = wage + income + profit + interest + indirect taxes = 480 + 160 + 120 + 50 + 40 = 850.